# Topological symmetry groups of the Generalized Petersen Graphs

### v1

@@ 8. The exceptional case $P(10,3)$
Recall from the proof of Proposition 8.7 that there is an order $2$ glide rotation $h$ of $S^{3}$ which interchanges $\Omega$ and $\Omega^{\prime}$, interchanging each vertex $i\in\Omega$ with the vertex $jklm\in\Omega^{\prime}$ where $i\not\in\{j,k,l,m\}$. Observe that $h$ interchanges $S$ and $S^{\prime}$ and leaves $\Gamma_{D4}$ setwise invariant inducing $\gamma_{3}$. Also, $\gamma_{3}$ commutes with $\gamma_{1}$ and $\gamma_{2}$, and hence $\langle\gamma_{1},\gamma_{2},\gamma_{3}\rangle=D_{4}\times\mathbb{Z}_{2}\leq\mathrm{TSG}(\Gamma_{D4})$. However, since $\mathrm{TSG}(\Gamma_{D4})$ is a proper subgroup of $\mathrm{TSG}(\Lambda{T})=S_{4}\times\mathbb{Z}_{2}$, and $D_{4}\times\mathbb{Z}_{2}$ is maximal in $S_{4}\times\mathbb{Z}_{2}$, it follows that $\mathrm{TSG}(\Gamma_{D4})=D_{4}\times\mathbb{Z}_{2}$.
Next, we replace a neighborhood of $u_{4}$ by one where each edge incident to $u_{4}$ is knotted around the next such edge as illustrated in Figure 6, except that now the chiral $3_{1}$ knot in the figure is replaced by the achiral $4_{1}$ knot. As explained in the proof of Proposition 6.5, this gives an order to the edges around $u_{4}$. We then apply the isometry group $\langle g,h\rangle$ to the neighborhood $N(u_{4})$, so that the edges in the orbit of $N(u_{4})$ also have an order. Since $g$ rotates $S_{1}$ and $S_{2}$, and $h$ is an involution interchanging $S_{1}$ and $S_{2}$, no vertex in the orbit of $u_{4}$ is fixed by a non-trivial element of $\langle g,h\rangle$. Thus, replacing the original neighborhoods of these vertices by the new ones yields a well-defined embedding $\Gamma_{Z4}$ of $P(10,3)$ such that $\mathrm{Z}_{4}\times\mathbb{Z}_{2}=\langle\gamma_{1},\gamma_{3}\rangle\leq\mathrm{TSG}(\Gamma_{Z4})$.
Now suppose that a homeomorphism $f^{\prime}$ of $(S^{3},\Gamma_{Z4})$ induces $\gamma_{2}$. Then $f^{\prime}$ fixes the edge $\overline{u_{5}u_{4}}$ and interchanges $\overline{u_{4}u_{3}}$ and $\overline{u_{4}v_{4}}$. But this means that $f^{\prime}$ reverses the order of the edges incident to $u_{4}$, which is impossible since each of these edges is knotted around the next one. Thus $\mathrm{TSG}(\Gamma_{Z4})$ is a proper subgroup of $\mathrm{TSG}(\Gamma_{D4})$. Since $\mathrm{Z}_{4}\times\mathbb{Z}_{2}$ is maximal in $D_{4}\times\mathbb{Z}_{2}$, we must have $\mathrm{TSG}(\Gamma_{Z4})=\mathrm{Z}_{4}\times\mathbb{Z}_{2}$.∎
@@ 9. Conclusion
In the table below we summarize our results for $P(n,k)$. A checkmark appears in the $\mathrm{TSG}(\Gamma)$ complete column if every subgroup of the automorphism group is realizable. A checkmark appears in the $\mathrm{TSG}_{+}(\Gamma)$ complete column if every subgroup of the automorphism group is positively realizable. We put an X in a column if some subgroups are not realizable or not positively realizable. The first three rows are the non-exceptional cases. The final five rows are all the exceptional cases except for $P(12,5)$ and $P(24,5)$, which will be analyzed in a subsequent paper.
@@ Acknowledgements
We thank AIM for its support of the REUF program which enabled us to start this project and to meet again the following year. We also thank ICERM for hosting us for our second meeting.

### v2

@@ 8. The exceptional case $P(10,3)$
Recall from the proof of Proposition 8.6 that there is an order $2$ glide rotation $h$ of $S^{3}$ which interchanges $\Omega$ and $\Omega^{\prime}$, interchanging each vertex $i\in\Omega$ with the vertex $jklm\in\Omega^{\prime}$ where $i\not\in\{j,k,l,m\}$. Observe that $h$ interchanges $S$ and $S^{\prime}$ and leaves $\Gamma_{D4}$ setwise invariant inducing $\gamma_{3}$. Also, $\gamma_{3}$ commutes with $\gamma_{1}$ and $\gamma_{2}$, and hence $\langle\gamma_{1},\gamma_{2},\gamma_{3}\rangle=D_{4}\times\mathbb{Z}_{2}\leq\mathrm{TSG}(\Gamma_{D4})$. However, since $\mathrm{TSG}(\Gamma_{D4})$ is a proper subgroup of $\mathrm{TSG}(\Lambda_{T})=S_{4}\times\mathbb{Z}_{2}$, and $D_{4}\times\mathbb{Z}_{2}$ is maximal in $S_{4}\times\mathbb{Z}_{2}$, it follows that $\mathrm{TSG}(\Gamma_{D4})=D_{4}\times\mathbb{Z}_{2}$.
Next, we replace a neighborhood of $u_{4}$ by one where each edge incident to $u_{4}$ is knotted around the next such edge as illustrated in Figure 7, except that now the chiral $3_{1}$ knot in the figure is replaced by the achiral $4_{1}$ knot. As explained in the proof of Proposition 6.5, this gives an order to the edges around $u_{4}$. We then apply the isometry group $\langle g,h\rangle$ to the neighborhood $N(u_{4})$, so that the edges in the orbit of $N(u_{4})$ also have an order. Since $g$ rotates $S_{1}$ and $S_{2}$, and $h$ is an involution interchanging $S_{1}$ and $S_{2}$, no vertex in the orbit of $u_{4}$ is fixed by a non-trivial element of $\langle g,h\rangle$. Thus, replacing the original neighborhoods of these vertices by the new ones yields a well-defined embedding $\Gamma_{Z4}$ of $P(10,3)$ such that $\mathbb{Z}_{4}\times\mathbb{Z}_{2}=\langle\gamma_{1},\gamma_{3}\rangle\leq\mathrm{TSG}(\Gamma_{Z4})$.
Now suppose that a homeomorphism $f^{\prime}$ of $(S^{3},\Gamma_{Z4})$ induces $\gamma_{2}$. Then $f^{\prime}$ fixes the edge $\overline{u_{5}u_{4}}$ and interchanges $\overline{u_{4}u_{3}}$ and $\overline{u_{4}v_{4}}$. But this means that $f^{\prime}$ reverses the order of the edges incident to $u_{4}$, which is impossible since each of these edges is knotted around the next one. Thus $\mathrm{TSG}(\Gamma_{Z4})$ is a proper subgroup of $\mathrm{TSG}(\Gamma_{D4})$. Since $\mathbb{Z}_{4}\times\mathbb{Z}_{2}$ is maximal in $D_{4}\times\mathbb{Z}_{2}$, we must have $\mathrm{TSG}(\Gamma_{Z4})=\mathbb{Z}_{4}\times\mathbb{Z}_{2}$.∎
@@ 9. Conclusion
In the table below we summarize our results for $P(n,k)$. A checkmark in the “$\mathrm{TSG}(\Gamma)$ complete” column means that every subgroup of the automorphism group is realizable. A checkmark $\checkmark$ in the “$\mathrm{TSG}_{+}(\Gamma)$ complete” column means that every subgroup of the automorphism group is positively realizable. An X means that some subgroups are not realizable or not positively realizable. The first three rows are in the non-exceptional cases. The final five rows are all the exceptional cases except for $P(12,5)$ and $P(24,5)$, which will be analyzed in a subsequent paper.
@@
Table 1. This table summarizes our results. Complete means that all of the subgroups are realized. The first three rows are in the non-exceptional cases
@@ Acknowledgements
We thank AIM for its support of the REUF program which enabled us to start this project and supported our meeting at the JMM. We also thank ICERM for hosting our second meeting. Finally, we thank the referee for helpful suggestions.